The length and width of a rectangle are given by consecutive integers.
The area of the rectangle is 110 m^2.
Find the length of the diagonal of the rectangle.

Found 2 solutions by richard1234, oberobic:
Answer by richard1234(7193)

It looks like you pressed "enter" after every word...if you're doing that please don't do that.

110 can be factored as 10*11. Hence the dimensions are 10 m, 11 m, and the length of the diagonal is .

Answer by oberobic(2304)

L = x
W = x+1
x(x+1) = 110
x^2 +x = 110
complete the square
.
x^2 +x + 1/4 = 110 + 1/4
(x + 1/2)^2 = sqrt(110+1/4)
.
(x + 1/2)^2 = (x+1/2)(x+1/2) = x^2 +1/2x + 1/2x + 1/4 = x^2 +x + 1/4
.
x + 1/2 = +sqrt(110 1/4) = 10.5
.
x = 10
.
x+1 = 11
.
Answer: The rectangle is 10 by 11 m.
.
Done.